day shift 07/22/99

John Shields (jshields@ksera)
Thu, 22 Jul 1999 15:36:16 -0500

Shift Summary for day shift, 07/22/99

On Shift: J. Shields, B.Hsiung

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 1.0

Accounting of lost time:

main injector power supply problems, P1 line position problem, and
Tevatron quenches.

========================================================================
Runs ended this shift:
Run 13974 - Pedestal run
Run 13975 - laser scan

Tapes brought to FCC this shift:
(including tapes for online split)

========================================================================
Accesses during shift:

========================================================================
Overview of shift:

(the blank space above pretty much sums it up)

the dehumidifier was worked on (and hopefully fixed), since it tripped
last night.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
07:00 beam off when shift crew arrives. Estimate is
for 8:30

09:00 control room lied. No beam yet. New estimate
is half an hour.

09:20 Went to Feynman Center to get a NEW DLT cleaning
tape. It worked much better than the old,
used-up ones lying around KTeV.
Bob notes that each cleaning tape should only
be used 20 times.

10:45 Still no beam, but Chan 13 has a new update!
Estimate is half an hour.

11:15 Although the above update implies that beam is
imnient, Eliot bravely risks a swap: put back
old AZLAZ since 2 new ones caused missing ffff word
at beginning of run. (missing bit 8)

12:45 Still no beam, but the intrepid Chan 13 has yet ANOTHER
update: Estimate is (say it *WITH* me now..) Half an hour.

12:50 swapped a keyboard with PREP.

13:52 Control room finally checks back into reality. latest
update: "Experts investigating a position problem in the
P1 line. *NO* estimate at this time."

14:15 Elizabeth calls from the airport and asks the intrepid
shift crew to e-mail some random dude at UCLA to let him
know that her flight was delayed. Bored out of their
minds, the shift crew was happy to oblige.

15:09 Yet another update from Chan 13. New estimate for
beam is -- what else??? -- half an hour.

15:18 Control Room called and said due to heat, etc.
Fermilab agreed to reduce the load on Edison by
shutting off beam tommorrow from around 9:30 am to
5 pm.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Patti and Rick K. both asked to be paged if beam ever comes back.